OpenStudy (anonymous):

-7(8x-3)+ x use distributive property to simplify the expression

Parth (parthkohli):

\[= \rm \left(-7\cdot 8x\right) - \left(-7 \cdot -3 \right)\]

OpenStudy (anonymous):

thanks